# Universal Transformers

### v1

@@ Appendix A Detailed Schema of the Universal Transformer
17 1.6 $\scriptstyle\pm$0.8 1.4 $\scriptstyle\pm$3.4 44.7 $\scriptstyle\pm$16.6 51.1 $\scriptstyle\pm$12.3
@@
avg 0.73 $\scriptstyle\pm$0.89 1.12 $\scriptstyle\pm$1.62 6.39 $\scriptstyle\pm$3.22 11.21 $\scriptstyle\pm$6.62
Appendix C bAbI Attention Visualization
We present visualization of the attention distributions on bAbI tasks for a couple of examples. The visualization of attention weights is over different time steps based on different heads over all the facts in the story and a question. Different color bars on the left side indicate attention weights based on different heads (4 heads in total).

### v2

@@ D.4 Learning to Execute (LTE)
17 1.6 $\scriptstyle\pm$0.8 1.4 $\scriptstyle\pm$3.4 43.7 $\scriptstyle\pm$18.6 51.1 $\scriptstyle\pm$12.9
@@
avg 0.73 $\scriptstyle\pm$0.89 1.12 $\scriptstyle\pm$1.62 6.34 $\scriptstyle\pm$3.32 11.21 $\scriptstyle\pm$6.65
Appendix F bAbI Attention Visualization
We present a visualization of the attention distributions on bAbI tasks for a couple of examples. The visualization of attention weights is over different time steps based on different heads over all the facts in the story and a question. Different color bars on the left side indicate attention weights based on different heads (4 heads in total).